# Can Vision Language Models Understand Mimed Actions?

### v1

@@ 1 Introduction
Figure 1: Simplified illustration of a sample in Mime shown with a single frame from a video of a 3D male character miming a basketball shot. Humans achieve almost perfect accuracy on identifying mimed actions regardless of evaluation format, adversarial perturbations, and the absence of any salient context (e.g., basketball, court, basketball uniform), while VLMs struggle without salient context.
@@ 2 Mime
2.1 Motion Capture
@@
2.2 Blender File Creation
@@ 4.3 Angle and Gender Variations
Model Eval Rotation Angle Avg. Std. $\downarrow$ 0°$\degree$ 90°$\degree$ 180°$\degree$ 270°$\degree$ Qwen 2.5 VL (3B) MC 34.9 34.9 32.6 32.6 33.7 1.2 FF 2.3 1.2 0.0 1.2 1.2 0.8 Qwen 2.5 VL (7B) MC 39.5 39.5 50.0 43.0 43.0 4.3 FF 5.8 7.0 3.5 8.1 6.1 1.7 Phi 3.5 MC 29.1 31.4 33.7 33.7 32.0 1.9 FF 2.3 5.8 3.5 3.5 3.8 1.3 InternVL2.5 (8B) MC 31.4 36.0 33.7 37.2 34.6 2.2 FF 2.3 7.0 7.0 4.7 5.2 1.9 GPT-4o Mini MC 41.9 47.7 43.0 47.7 45.1 2.6 FF 11.6 15.1 13.9 13.9 13.7 1.3 Gemini 1.5 Flash MC 52.3 47.7 52.3 53.5 51.5 2.2 FF 19.8 18.6 17.4 23.3 19.8 2.2 Human MC 99.6 98.8 98.8 98.7 99.0 0.4 FF 89.5 95.0 90.7 85.1 90.1 3.5
@@
Table 3: Performance comparison on Mime for gender variations. $\Delta$ is shown in blue if ♂--$\geq 0$ and in orange otherwise. Similar to angle variation results, results for VLMs vary largely depending on the gender without a consistent performance advantage of a certain gender, while human performance is consistent.
@@ 5.3 Failure Mode Analysis
We find that, with the blank background, Gemini 1.5 Flash generates a description of the shown mimed action that is only partially correct 54% of the time and completely incorrect 16% of the time (e.g., They wind up their arm as if holding a ball, then perform a throwing motion with their arm and hand extending forward for arm curls). In 13% of instances, the description is correct, but it is interpreted incorrectly, leading to an incorrect prediction (e.g., predicts bowling a ball after generating They start with a wind-up motion, bringing their arm back, then swing forward as if releasing a ball for baseball pitch). When shown an aligned background, 43% of predictions that were incorrect with the blank background become correct predictions. With the misaligned background, 24% of all predictions are confused by irrelevant context provided by the background (e.g., predicts conducting an orchestra for climbing given a concert hall background), which leads to a drop in proportion of predictions that had completely or partially correct descriptions of the shown mimed action.
We share the full statistics and more examples of each mode of failure in Appendix  E. Other VLMs also show similar failure patterns, with most instances of failure caused by partial or completely incorrect descriptions of the mimed actions. Overall, these results indicate that future research should prioritize training VLMs that can accurately describe the human gestures they observe.
@@ 6.2 Action Recognition
While mimed action understanding is an instrumental step towards general NVC understanding, it is also highly related to action understanding. Many datasets exist for evaluating whether machine learning models understand human actions Kong and Fu (2022); Sun et al. (2022), with early work focused on sporting actions Kuehne et al. (2011); Soomro et al. (2012); Karpathy et al. (2014); Idrees et al. (2017) and recent work expanding to a larger scope and scale, including daily activities that are crowdsourced Sigurdsson et al. (2016); Damen et al. (2018) and extracted from YouTube Heilbron et al. (2015); Xu et al. (2016); Krishna et al. (2017); Sanabria et al. (2018); Zhou et al. (2018); Miech et al. (2019), Tumblr Li et al. (2016), Flickr Anne Hendricks et al. (2017), or movies Torabi et al. (2015); Rohrbach et al. (2015, 2017).  Miech et al. (2020) takes a step further to evaluate action recognition robustness with a dataset that contains rare activities (e.g., blending phone and cutting keyboard). While these datasets collectively cover more than hundreds of different classes, none of them contain mimed actions. As such, even if VLMs perform well on these datasets, it is unclear whether they have a robust understanding of complex human body motions or if they are relying on spurious correlations provided by the salient context.
The CMU-MMAC Database la Torre et al. (2008) is the most similar to Mime in that it provides animated videos of motion capture data. However, its videos are extremely simplified, featuring a black background with yellow stick figures, and do not include finger and thumb joints, which lowers fidelity to the captured actions. IEMOCAP Busso et al. (2008) is another dataset based on motion capture data but it is focused on emotion prediction in dyadic conversations and motion capture is only collected for the face, head, and hands. van Nispen et al. (2017) is also similar in that it presents a dataset of footage of human participants pantomiming various objects, but it is limited to hand gestures and its primary purpose is studying human psychology by analyzing how humans vary in how they mime various tools.
@@ Acknowledgment
Next, Mime only consists of animations that are based on motion capture data from two actors. However, one of these actors is an Asian male non-professional actor and the other is an European female professional actor. Despite minimal overlap in demographics of these two actors, our human evaluation results show that there is not a higher recognition accuracy for samples that come from one actor over those of another.10101099.0% vs. 98.8% in multiple choice format and 92.6% vs. 90.6% for free-form format. A t-test indicates insignificant differences in these accuracies at $p<0.01$ ($p=0.71$ and $p=0.03$, respectively). In other words, there may be differences in how people enact actions as mimes, but they are not significant enough to affect recognition, at least for the pool of 60 participants that we recruited. We believe these results should address concerns of representation bias of the mimed actions in Mime.
@@ Appendix C Fine-tuning Details
[…]
@@ Appendix D Human Evaluation Details
Without any prior guidance, each participant is asked to answer the question in free-form format first after watching a sample in Mime and then answer the same question with the multiple choice format. The interface shown to our participants is illustrated in Figure 6. While this setup is efficient for collecting both free-form and multiple choice format results from a single participant, options shown for the multiple choice format in prior samples may provide contextual information for free-form format answers in the remaining samples. However, we find this effect to be negligible in comparison to the large performance gap between humans and models: human performance with the free-form format on the first samples that they annotate is ~88% while that of the remaining samples is ~93%. Each participant annotates half of the total samples (43) in one of the configurations shown in Figure 3 in order to keep annotator burden low.
@@ Appendix E Failure Mode Anlaysis Details
Table 6: Gemini 1.5 Flash with CoT’s prediction category distributions (*Note % indicates proportion of predictions, including correct ones) and corresponding examples, as discussed in §5.3. ○: completely incorrect description. ◑: partially correct description of action. ●: correct description with only minor discrepancies but incorrect interpretation. ✓: correct description and valid alternative interpretation. [Uncaptioned image]: predictions helped by context (background is aligned with mimed action). This is the only row where the proportion is not of failures but of accurately predicted instances. [Uncaptioned image]: predictions confused by context (background is adversarial). Correct descriptions are in blue and incorrect descriptions are in orange.

### v2

@@ 1 Introduction
Figure 1: Simplified illustration of a sample in Mime shown with a few frames from a video of a 3D male character miming a basketball shot in a living room. Humans achieve almost perfect accuracy on identifying mimed actions regardless of evaluation format, adversarial perturbations, and the absence of salient context (e.g., basketball, court, basketball outfit), while VLMs struggle without salient context.
@@ 2 Mime
2.1 Collecting Motion Capture Data
@@
2.2 Creating Blender Files
@@ 4.3 Angle and Gender Variations
[…]
@@
Table 3: Performance comparison on Mime for gender variations. $\Delta$ is shown in blue if $\geq 0$ and in orange otherwise. Similar to angle variation results, results for VLMs vary largely depending on the gender without a consistent performance advantage of a certain gender, while human performance is consistent.
@@ 5.3 Failure Mode Analysis
We find that, with the blank background, Gemini 1.5 Flash generates a description of the shown mimed action that is only partially correct 54% of the time and completely incorrect 16% of the time (e.g., They wind up their arm as if holding a ball, then perform a throwing motion with their arm and hand extending forward for arm curls). In 13% of instances, the description is correct, but it is interpreted incorrectly, leading to an incorrect prediction (e.g., predicts bowling a ball after generating They start with a wind-up motion, bringing their arm back, then swing forward as if releasing a ball for baseball pitch). When shown an aligned background, 43% of predictions that were incorrect with the blank background become correct predictions. With the misaligned background, 24% of all predictions are confused by irrelevant context provided by the background (e.g., predicts conducting an orchestra for climbing given a concert hall background), which leads to a drop in proportion of predictions that had completely or partially correct descriptions of the shown mimed action.101010 We share the full statistics and more examples of each mode of failure in Appendix  E. Other VLMs also show similar failure patterns, with most instances of failure caused by partial or completely incorrect descriptions of the mimed actions. Overall, these results indicate that future research should prioritize training VLMs that can accurately describe the human gestures they observe.
@@ 6.2 Action Recognition
While mimed action understanding is an instrumental step towards general NVC understanding, it is also highly related to action understanding. Many datasets exist for evaluating whether machine learning models understand human actions Kong and Fu (2022); Sun et al. (2022), with early work focused on sporting actions Kuehne et al. (2011); Soomro et al. (2012); Karpathy et al. (2014); Idrees et al. (2017) and recent work expanding to a larger scope and scale, including daily activities that are crowdsourced Sigurdsson et al. (2016); Damen et al. (2018) and extracted from YouTube Heilbron et al. (2015); Xu et al. (2016); Krishna et al. (2017); Kay et al. (2017); Sanabria et al. (2018); Zhou et al. (2018); Miech et al. (2019); Weinzaepfel and Rogez (2021), Tumblr Li et al. (2016), Flickr Anne Hendricks et al. (2017), or movies Torabi et al. (2015); Rohrbach et al. (2015, 2017).  Miech et al. (2020) takes a step further to evaluate action recognition robustness with a dataset that contains rare activities (e.g., blending phone and cutting keyboard). While these datasets collectively cover more than hundreds of different classes, none of them contain mimed actions. As such, even if VLMs perform well on these datasets, it is unclear whether they have a robust understanding of complex human body motions or if they are relying on spurious correlations provided by the salient context.
The CMU-MMAC Database la Torre et al. (2008) is similar to Mime in that it contains animated videos of motion capture data, but the animations are extremely simple, featuring a black background with yellow stick figures, and do not include finger and thumb joints, which lowers fidelity to the captured actions. IEMOCAP Busso et al. (2008) is also based on motion capture data but it is focused on emotion prediction in dyadic conversations and motion capture is only collected for the face, head, and hands. van Nispen et al. (2017) presents a dataset of footage of human participants pantomiming various objects, but it is limited to hand gestures. Lastly, the Mimetics dataset Weinzaepfel and Rogez (2021) is the most similar to Mime in that in contains live action footage videos of mimed actions extracted from YouTube with varying amounts of relevant context provided. However, it lacks the flexibility to systematically adjust the amount of relevant context provided in each video with precision for conducting the ablative analysis possible with Mime. Moreover, the Mimetics dataset may be included in the training of VLMs as the videos are from YouTube, and therefore Mime serves as a more reliable benchmark for mimed action understanding unaffected by data leakage.
@@ Acknowledgment
Next, Mime only consists of animations that are based on motion capture data from two actors. However, one of these actors is an Asian male non-professional actor and the other is an European female professional actor. Despite minimal overlap in demographics of these two actors, our human evaluation results show that there is not a higher recognition accuracy for samples that come from one actor over those of another.11111199.0% vs. 98.8% in multiple choice format and 92.6% vs. 90.6% for free-form format. A t-test indicates insignificant differences in these accuracies at $p<0.01$ ($p=0.71$ and $p=0.03$, respectively). In other words, there may be differences in how people enact actions as mimes, but they are not significant enough to affect recognition, at least for the pool of 60 participants that we recruited. We believe these results should address concerns of representation bias of the mimed actions in Mime.
@@ Appendix C Fine-tuning Details
The $n=5$ folds that we use for N-fold training for fine-tuning experiments are shown in Table 5. During FT, only the vision encoder is trained, while the text encoder remains frozen. We train for 7 epochs with an initial learning rate of 2ee-5, following a cosine learning rate schedule. The batch size is set to 8, and we use the AdamW optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. To optimize the balance between computational speed and precision, BF16 and TF32 are enabled. All models are trained using 2× A100 GPUs.
@@ Appendix D Human Evaluation Details
Without any prior guidance, each participant is asked to answer the question in free-form format first after watching a sample in Mime and then answer the same question with the multiple choice format. The interface shown to our participants is illustrated in Figure 6. While this setup is efficient for collecting both free-form and multiple choice format results from a single participant, options shown for the multiple choice format in prior samples may provide contextual information for free-form format answers in the remaining samples. However, we find this effect to be negligible in comparison to the large performance gap between humans and models: human performance with the free-form format on the first samples that they annotate is ˜88% while that of the remaining samples is ˜93%. Each participant annotates half of the total samples (43) in one of the configurations shown in Figure 3 in order to keep annotator burden low.
@@ Appendix E Failure Mode Anlaysis Details
Table 6: Gemini 1.5 Flash with CoT’s prediction category distributions (*Note % indicates proportion of predictions, including correct ones) and corresponding examples, as discussed in §5.3. ○: completely incorrect description. ◑: partially correct description of action. ●: correct description with only minor discrepancies but incorrect interpretation. ✓: correct description and valid alternative interpretation. Correct descriptions are in blue and incorrect descriptions are in orange.